# Epimorphisms between finitely generated algebras

### v2

@@ 1. Introduction
The interest of the weak ES property is twofold. On the one hand, most quasivarieties in which epimorphisms need not be surjective lack the weak ES property as well. For instance, the inclusion map of $\mathbb{Z}$ into $\mathbb{Z}[1/p]$ for a prime $p$ is a nonsurjective epimorphism between finitely generated rings, whence the variety of rings lacks the weak ES property. On the other hand, from a logical standpoint, the weak ES property is the algebraic counterpart of the Beth definability property, i.e., the demand that all implicit definitions can be made explicit (see [17, Thm. 5.6.10] and [5, Thm. 3.14]). In essence, the latter means that if an element of a structure satisfying some property is unique when it exists, then it is definable by a term of the language. For instance, the failure of the weak ES property in rings amounts to the fact that multiplicative inverses are implicitly, but not explicitly, definable.
@@
We show that if $\mathsf{K}$ has a near unanimity term of arity $n$, we may assume that the algebra $\bm{B}$ is a subdirect product $\bm{B}\leqslant\bm{C}_{1}\times\dots\times\bm{C}_{n-1}$ of relatively finitely subdirectly irreducible factors (Theorem 4.3). Furthermore, we prove that if $\mathsf{K}$ is a congruence permutable variety, we may assume that the algebra $\bm{B}$ is finitely subdirectly irreducible (Theorem 5.3). Notably, the first result applies to every variety with a lattice reduct and the second to every variety with a group reduct (see, e.g., [7, p. 79 & Thm. II.12.2]), thus covering most examples from algebra and logic. [^1] We remark that Theorems 4.3 and 5.3 are slightly stronger than suggested here, but their formulation would require additional definitions.
@@
We remark that Theorems 4.3 and 5.3 strengthen similar observations on the surjectivity of all $\mathsf{K}$-epimorphisms [8, Thms. 18 & 22] (here Theorems 4.2 and 5.2). Both our improvements require new proof strategies. For instance, [8, Thm. 22] states that all $\mathsf{K}$-epimorphisms are surjective for an arithmetical variety $\mathsf{K}$ whose class of finitely subdirectly irreducible members is universal provided that no proper inclusion map $f\colon\bm{A}\to\bm{B}$, where $\bm{B}$ is finitely subdirectly irreducible, is a $\mathsf{K}$-epimorphism. We show that, for the weak ES property, this is true under the sole assumption of congruence permutablity (Theorem 5.3). This requires to dispense with the main ingredients of [8], namely, the theory of definability in [10], sheaf representations [16, 23], and the infinitary Baker-Pixley Theorem [31, 11]. Instead, we capitalize on the new notion of a full subalgebra and its interaction with subdirect representations.
@@ 2. Varieties and quasivarieties
In view of Birkhoff’s and Maltsev’s Theorems, varieties and quasivarieties coincide with the classes of algebras axiomatized by equations and quasiequations, respectively (see, e.g., [7, Thms. II.11.9 & V.2.25]). While every variety is a quasivariety, the converse is not true in general. We denote the least variety and the least quasivariety containing a class of algebras $\mathsf{K}$ by $\mathbb{V}(\mathsf{K})$ and $\mathbb{Q}(\mathsf{K})$, respectively. A variety (resp. quasivariety) $\mathsf{K}$ is finitely generated when $\mathsf{K}=\mathbb{V}(\mathsf{M})$ (resp. $\mathsf{K}=\mathbb{Q}(\mathsf{M})$) for a finite set of finite algebras $\mathsf{M}$.
@@
Let $\mathsf{K}$ be a quasivariety. An algebra $\bm{A}\in\mathsf{K}$ is said to be subdirectly irreducible relative to $\mathsf{K}$ (RSI, for short) when for every subdirect embedding $f\colon\bm{A}\to\prod_{i\in I}\bm{B}_{i}$ with $\{\bm{B}_{i}:i\in I\}\subseteq\mathsf{K}$ there exists $i\in I$ such that $p_{i}\circ f\colon\bm{A}\to\bm{B}_{i}$ is an isomorphism. In case this happens whenever the index set $I$ is finite, we say that $\bm{A}$ is finitely subdirectly irreducible relative to $\mathsf{K}$ (RFSI, for short). The classes of RSI and RFSI members of $\mathsf{K}$ will be denoted by $\mathsf{K}_{\textsc{rsi}}$ and $\mathsf{K}_{\textsc{rfsi}}$, respectively. When $\mathsf{K}$ is a variety, the requirement that $\{\bm{B}_{i}:i\in I\}$ is a subset of $\mathsf{K}$ in the above definitions can be harmlessly dropped and we simply say that $\bm{A}$ is subdirectly irreducible (SI) or finitely subdirectly irreducible (FSI) (i.e., we drop the “relative to $\mathsf{K}$”). In this case, we also write $\mathsf{K}_{\textsc{si}}$ and $\mathsf{K}_{\textsc{fsi}}$ instead of $\mathsf{K}_{\textsc{rsi}}$ and $\mathsf{K}_{\textsc{rfsi}}$.
@@
Notably, the RSI and RFSI members of a quasivariety $\mathsf{K}$ can be recognized by looking at the structure of their lattices of $\mathsf{K}$-congruences. More precisely, given $\bm{A}\in\mathsf{K}$, let
@@
One of the main consequences of congruence distributivity is the following (see, e.g., [3, Lem. 5.9]): [^2] Although we will not need it, Jónsson’s Lemma can be generalized to congruence distributive quasivarieties [30, Thm. 1.7].
Jónsson’s Lemma 2.6.
@@ 3. Epimorphism surjectivity
Suppose first that $\bm{A}\leqslant\bm{B}$ is not epic in $\mathsf{K}$. Therefore, there exist $\bm{C}\in\mathsf{K}$ and two homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g\neq h$. As $g\neq h$ there exists also $b\in B$ such that $g(b)\neq h(b)$. By the Subdirect Decomposition Theorem 2.3 there exists $\bm{D}\in\mathsf{K}_{\textsc{rsi}}$ and a homomorphism $f\colon\bm{C}\to\bm{D}$ such that $f(g(b))\neq f(h(b))$. Thus, $f\circ g$ and $f\circ h$ differ, but their restrictions to $A$ coincide. Consequently, we may assume that $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$ (otherwise, we replace $\bm{C}$ by $\bm{D}$ and $h$ and $g$ by their compositions with $f$).
@@
Now, we turn to prove the converse. If condition (ii) holds, then it is clear that $\bm{A}\leqslant\bm{B}$ is not epic in $\mathsf{K}$. Then suppose that condition (i) holds. In this case, there are two distinct $\theta,\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ such that $\theta{\upharpoonright}_{A}=\phi{\upharpoonright}_{A}$. We first show that we may assume that $\theta=\textup{id}_{B}$. As $\theta\neq\phi$, we have $\theta\nsubseteq\phi$ or $\phi\nsubseteq\theta$. By symmetry, may assume that $\phi\nsubseteq\theta$, and hence that $\phi\neq\theta\cap\phi$. Since $(\theta\cap\phi){\upharpoonright}_{A}=\theta{\upharpoonright}_{A}\cap\phi{\upharpoonright}_{A}=\phi{\upharpoonright}_{A}$, Lemma 3.9 yields that $\theta\cap\phi=\textup{id}_{B}$. Thus, replacing $\theta$ with $\theta\cap\phi$ allows us to assume that $\theta=\textup{id}_{B}$, $\phi\neq\textup{id}_{B}$, and $\phi{\upharpoonright}_{A}=(\textup{id}_{B}){\upharpoonright}_{A}$.
@@
Let $\mathsf{K}$ be a quasivariety. If $\bm{B}\in\mathsf{K}$ and $\bm{A}\leqslant\bm{B}$ is full and not epic in $\mathsf{K}$, then $\bm{A}\in\mathbb{S}(\mathsf{K}_{\textup{{rfsi}}})$.
@@
$\textup{id}_{A}=\textup{id}_{B}{\upharpoonright}_{A}=(\mathsf{Ker}(g)\cap\mathsf{Ker}(h)){\upharpoonright}_{A}=\mathsf{Ker}(g){\upharpoonright}_{A}\cap\mathsf{Ker}(h){\upharpoonright}_{A}=\mathsf{Ker}(g){\upharpoonright}_{A}.$
Therefore, $g{\upharpoonright}_{A}$ is injective as desired. As a consequence, $g{\upharpoonright}_{A}\colon\bm{A}\to\bm{C}$ is an embedding. Since $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$, we conclude that $\bm{A}\in\mathbb{I}\mathbb{S}(\mathsf{K}_{\textup{{rfsi}}})\subseteq\mathbb{S}(\mathsf{K}_{\textup{{rfsi}}})$. ∎
@@
Corollary 3.11 cannot be strengthened by concluding that also $\bm{B}\in\mathbb{S}(\mathsf{K}_{\textup{{rfsi}}})$. To prove this, let $\bm{B}$ be the semilattice with meet-order $10<a<$ 1 and $\bm{A}$ its subalgebra with universe $\{0,1\}$. While $\bm{A}\leqslant\bm{B}$ is easily seen to be full in $\mathsf{SL}$, it is not epic because $\mathsf{SL}$ has the weak ES property (see, e.g., [21, p. 99]). Nonetheless, $\bm{B}$ is not a subalgebra of an FSI semilattice because, up to isorphism, the only FSI semilattice is $\bm{A}$. ∎
@@ 4. Quasivarieties with a near unanimity term
In the presence of a near unanimity term, the task of determining whether a quasivariety has the ES property is simplified by the following result [8, Thm. 18]: [^5] In the statement of [8, Thm. 18], the occurrences of $n-1$ in Theorem 4.2 are replaced by $n$. However, its proof yields the stronger version reported here (see [25]).
@@
ordered under the relation given by $\langle\phi_{1},\dots,\phi_{m-1}\rangle\leqslant\langle\eta_{1},\dots,\eta_{m-1}\rangle$ if and only if $\phi_{i}\subseteq\eta_{i}$ for every $i\leqslant m-1$. We will apply Zorn’s Lemma to obtain a maximal element of $X$. Clearly, $X$ contains $\langle\theta_{1},\dots,\theta_{m-1}\rangle$. Consider a nonempty chain $C$ in $X$. For each $i\leqslant m-1$ let $C_{i}$ be the projection of $C$ on the $i$-th coordinate. Observe that $C_{i}$ is a nonempty chain in $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ because $C$ is a nonempty chain in $X$. We will prove that $\langle\bigcup C_{1},\dots,\bigcup C_{m-1}\rangle$ is an upper bound of $C$ in $X$. Proposition 3.6 implies that each $\bigcup C_{i}$ is a $\mathsf{K}$-congruence of $\bm{A}$. Furthermore, as $\theta_{i}$ is contained in every member of $C_{i}$ and $C_{i}$ is nonempty, we have $\theta_{i}\subseteq\bigcup C_{i}$. Lastly, observe that
$\left(\bigcup C_{1}\right)\cap\dots\cap\left(\bigcup C_{m-1}\right)=\bigcup\{\phi_{1}\cap\dots\cap\phi_{m-1}:\phi_{i}\in C_{i}\text{ for }i\leqslant m-1\}=\theta_{1}\cap\dots\cap\theta_{m-1},$
@@
where the the first inclusion holds because $\phi_{i}\in C_{i}$ for every $i\leqslant m-1$, and the last equality holds because $\langle\phi_{1}^{\prime},\dots,\phi_{m-1}^{\prime}\rangle\in C$. Hence, $\phi_{1}\cap\dots\cap\phi_{m-1}=\theta$ because $\theta=\theta_{1}\cap\dots\cap\theta_{m-1}$. This establishes the above display and shows that $\langle\bigcup C_{1},\dots,\bigcup C_{m-1}\rangle$ is an upper bound of $C$ in $X$. By Zorn’s Lemma the poset $X$ has a maximal element $\langle\phi_{1},\dots,\phi_{m-1}\rangle$. In particular, $\theta=\phi_{1}\cap\dots\cap\phi_{m-1}$.
@@
As $\theta$ is $m$-irreducible and $\theta\neq\phi_{1}\cap\dots\cap\phi_{i-1}\cap\phi_{i+1}\cap\dots\cap\phi_{m-1}$, we obtain
@@
$\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),b\rangle\in\theta_{j}$
for every $j\leqslant n$. To this end, consider $j\leqslant n$. As $\langle a_{i},b\rangle\in\phi_{i}\subseteq\theta_{j}$ for every $i\leqslant n$ such that $i\neq j$, we obtain $\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),\varphi^{\bm{B}}(b,\dots,b,a_{j},b,\dots,b)\rangle\in\theta_{j}$. Furthermore, since $\varphi$ is a near unanimity term, we have $\varphi^{\bm{B}}(b,\dots,b,a_{j},b,\dots,b)=b$. Hence, $\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),b\rangle\in\theta_{j}$. This establishes the above display. Together with the assumption that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n}$, this implies $b=\varphi^{\bm{B}}(a_{1},\dots,a_{n})$. As $a_{1},\dots,a_{n}\in A$ and $\bm{A}\leqslant\bm{B}$, we conclude that $b\in A$, which is false. Hence, $\text{id}_{B}$ is $n$-irreducible as desired.
@@
The implication from left to right holds by Proposition 3.2. We will prove the other implication by contraposition. Suppose that $\mathsf{K}$ lacks the weak ES property. By Corollary 3.8 there exist $\bm{B}\in\mathsf{K}$ finitely generated and $\bm{A}\leqslant\bm{B}$ fully epic in $\mathsf{K}$. In view of Proposition 4.6, we may assume that $\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n}$ is a subdirect product for some $\bm{B}_{1},\dots,\bm{B}_{n}\in\mathsf{K}_{\textsc{rfsi}}$. ∎
@@
Let $\mathsf{K}$ be a class of lattices of bounded length such that $\mathbb{V}(\mathsf{K})$ is nontrivial. By Jónsson’s Lemma 2.6 the FSI members of $\mathbb{V}(\mathsf{K})$ belong to $\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})$. Since the members of $\mathsf{K}$ have length $\leqslant n$ for some $n\in\mathbb{Z}^{+}$, so do those of $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})$ by Łoś’ Theorem [7, Thm. V.2.9]. As the class operators $\mathbb{H}$ and $\mathbb{S}$ do not increase the cardinality of chains, we conclude that the FSI members of $\mathbb{V}(\mathsf{K})$ have length $\leqslant n$. Now, as $\mathbb{V}(\mathsf{K})$ is nontrivial, it has at least one SI member. Therefore, there exists $\bm{A}_{1}\in\mathbb{V}(\mathsf{K})_{\textsc{si}}$ of maximal length. Let $\textup{Cg}(a,b)$ be the monolith of $\bm{A}$. As $a\neq b$, we may assume that $a<b$. Moreover, let $\bm{A}_{2}$ be the two-element chain $0<1$ viewed as a lattice, which is also SI. Consider the subalgebra $\bm{A}$ of $\bm{A}_{1}\times\bm{A}_{2}$ with universe $(A_{1}\times\{0\})\cup({\uparrow}b\times\{1\})$, where ${\uparrow}b\coloneqq\{c\in A_{1}:b\leqslant c\}$, and let $\bm{B}\coloneqq\textup{Sg}^{\bm{A}_{1}\times\bm{A}_{2}}(A\cup\{\langle a,1\rangle\})$. Clearly, $\bm{B}\leqslant\bm{A}_{1}\times\bm{A}_{2}$ is a subdirect product. We will prove that $\bm{A}\leqslant\bm{B}$ is fully epic in $\mathbb{V}(\mathsf{K})$, thus showing that $\mathbb{V}(\mathsf{K})$ lacks the weak ES property.

### v3

@@ 1. Introduction
The interest of the weak ES property is twofold. On the one hand, it is often the case that quasivarieties in which epimorphisms need not be surjective lack the weak ES property as well. For instance, the inclusion map of $\mathbb{Z}$ into $\mathbb{Z}[1/p]$ for a prime $p$ is a nonsurjective epimorphism between finitely generated rings, whence the variety of rings lacks the weak ES property. More in general, the ES property and its weak version coincide for quasivarieties with the amalgamation property (see, e.g., [4, Thm. 1.3]). On the other hand, from a logical standpoint, the weak ES property is the algebraic counterpart of the (finite) Beth definability property, i.e., the demand that all implicit definitions can be made explicit (see [17, Thm. 5.6.10] and [5, Thm. 3.14]). In essence, the latter means that if an element of a structure satisfying some property is unique when it exists, then it is definable by a term of the language. For instance, the failure of the weak ES property in rings amounts to the fact that multiplicative inverses are implicitly, but not explicitly, definable.
@@
We show that if $\mathsf{K}$ has a near unanimity term of arity $n$, we may assume that the algebra $\bm{B}$ is a subdirect product $\bm{B}\leqslant\bm{C}_{1}\times\dots\times\bm{C}_{n-1}$ of relatively finitely subdirectly irreducible factors (Theorem 4.3). Furthermore, we prove that if $\mathsf{K}$ is a congruence permutable variety, we may assume that the algebra $\bm{B}$ is finitely subdirectly irreducible (Theorem 5.3). Notably, the first result applies to every variety of algebras with a lattice reduct and the second to every variety of algebras with a group reduct (see, e.g., [7, p. 79 & Thm. II.12.2]), thus covering most examples from algebra and logic. [^1] We remark that Theorems 4.3 and 5.3 are slightly stronger than suggested here, but their formulation would require additional definitions.
@@
We remark that Theorems 4.3 and 5.3 strengthen similar observations on the surjectivity of all $\mathsf{K}$-epimorphisms [8, Thms. 6.4 & 6.8] (here Theorems 4.2 and 5.2). Both our improvements require new proof strategies. For instance, [8, Thm. 6.8] states that all $\mathsf{K}$-epimorphisms are surjective for an arithmetical variety $\mathsf{K}$ whose class of finitely subdirectly irreducible members is universal provided that no proper inclusion map $f\colon\bm{A}\to\bm{B}$, where $\bm{B}$ is finitely subdirectly irreducible, is a $\mathsf{K}$-epimorphism. We show that, for the weak ES property, this is true under the sole assumption of congruence permutability (Theorem 5.3). This allows us to dispense with the main ingredients of [8], namely, the theory of definability in [10], sheaf representations [16, 23], and the infinitary Baker-Pixley Theorem [30] (see also [11]). Instead, we capitalize on the new notion of a full subalgebra and its interaction with subdirect representations.
@@ 2. Varieties and quasivarieties
In view of Birkhoff’s and Maltsev’s Theorems, varieties and quasivarieties coincide with the classes of algebras axiomatized by equations and quasiequations, respectively (see, e.g., [7, Thms. II.11.9 & V.2.25]). While every variety is a quasivariety, the converse is not true in general. We denote the least variety and the least quasivariety containing a class of algebras $\mathsf{K}$ by $\mathbb{V}(\mathsf{K})$ and $\mathbb{Q}(\mathsf{K})$, respectively. A variety (resp. quasivariety) $\mathsf{K}$ is finitely generated when $\mathsf{K}=\mathbb{V}(\mathsf{M})$ (resp. $\mathsf{K}=\mathbb{Q}(\mathsf{M})$) for a finite set $\mathsf{M}$ of finite algebras.
@@
Let $\mathsf{K}$ be a quasivariety. An algebra $\bm{A}\in\mathsf{K}$ is said to be subdirectly irreducible relative to $\mathsf{K}$ (RSI, for short) when for every subdirect embedding $f\colon\bm{A}\to\prod_{i\in I}\bm{B}_{i}$ with $\{\bm{B}_{i}:i\in I\}\subseteq\mathsf{K}$ there exists $i\in I$ such that $p_{i}\circ f\colon\bm{A}\to\bm{B}_{i}$ is an isomorphism. In case this happens whenever the index set $I$ is finite, we say that $\bm{A}$ is finitely subdirectly irreducible relative to $\mathsf{K}$ (RFSI, for short).222 When the language under consideration is clear, we adopt the following convention: the direct product of an empty family of algebras is the trivial algebra in the language under consideration. Consequently, trivial algebras are considered neither RSI nor RFSI. The classes of RSI and RFSI members of $\mathsf{K}$ will be denoted by $\mathsf{K}_{\textsc{rsi}}$ and $\mathsf{K}_{\textsc{rfsi}}$, respectively. When $\mathsf{K}$ is a variety, the requirement that $\{\bm{B}_{i}:i\in I\}$ is a subset of $\mathsf{K}$ in the above definitions can be harmlessly dropped and we simply say that $\bm{A}$ is subdirectly irreducible (SI) or finitely subdirectly irreducible (FSI) (i.e., we drop the “relative to $\mathsf{K}$”). In this case, we also write $\mathsf{K}_{\textsc{si}}$ and $\mathsf{K}_{\textsc{fsi}}$ instead of $\mathsf{K}_{\textsc{rsi}}$ and $\mathsf{K}_{\textsc{rfsi}}$.
@@
Notably, the RSI and RFSI members of a quasivariety $\mathsf{K}$ can be recognized by looking at the structure of their lattices of $\mathsf{K}$-congruences. To explain how, we recall that an element $a$ of a complete lattice $\bm{A}$ is
  1. (i)
    completely meet irreducible if for every family $\{a_{i}:i\in I\}\subseteq A$,
    $a=\bigwedge_{i\in I}a_{i}\textrm{ implies }a=a_{i}\textrm{ for some }i\in I;$
  2. (ii)
    meet irreducible if the above display holds whenever $I$ is finite.
As the set of indices $I$ is allowed to be empty, the greatest element of $\bm{A}$ is never meet irreducible.
Given $\bm{A}\in\mathsf{K}$, let
@@
The following theorem can be derived from Jónsson’s Lemma (see, e.g., [3, Lem. 5.9]), which is one of the main consequences of congruence distributivity.
Theorem 2.6.
@@ 3. Epimorphism surjectivity
Suppose first that $\bm{A}\leqslant\bm{B}$ is not epic in $\mathsf{K}$. Therefore, there exist $\bm{C}\in\mathsf{K}$ and two homomorphisms $g,h\colon\bm{B}\to\bm{C}$ such that $g{\upharpoonright}_{A}=h{\upharpoonright}_{A}$ and $g\neq h$. As $g\neq h$ there exists also $b\in B$ such that $g(b)\neq h(b)$. By the Subdirect Decomposition Theorem 2.3 there exist $\bm{D}\in\mathsf{K}_{\textsc{rsi}}$ and a homomorphism $f\colon\bm{C}\to\bm{D}$ such that $f(g(b))\neq f(h(b))$. Thus, $f\circ g$ and $f\circ h$ differ, but their restrictions to $A$ coincide. Consequently, we may assume that $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$ (otherwise, we replace $\bm{C}$ by $\bm{D}$ and $h$ and $g$ by their compositions with $f$).
@@
Now, we prove the converse. If condition (ii) holds, then it is clear that $\bm{A}\leqslant\bm{B}$ is not epic in $\mathsf{K}$. Then suppose that condition (i) holds. In this case, there are two distinct $\theta,\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ such that $\theta{\upharpoonright}_{A}=\phi{\upharpoonright}_{A}$. Then Lemma 3.9 yields $\theta\cap\phi=\textup{id}_{B}$. As $\theta\neq\phi$, we have $\theta\nsubseteq\phi$ or $\phi\nsubseteq\theta$. By symmetry, we may assume that $\phi\nsubseteq\theta$, and hence that $\phi\neq\theta\cap\phi$. From $\phi{\upharpoonright}_{A}=\theta{\upharpoonright}_{A}$ it follows that $\phi{\upharpoonright}_{A}=\theta{\upharpoonright}_{A}\cap\phi{\upharpoonright}_{A}=(\theta\cap\phi){\upharpoonright}_{A}$. Thus, replacing $\theta$ with $\theta\cap\phi$ allows us to assume that $\theta=\textup{id}_{B}$, $\phi\neq\textup{id}_{B}$, and $\phi{\upharpoonright}_{A}=(\textup{id}_{B}){\upharpoonright}_{A}$.
@@
Let $\mathsf{K}$ be a quasivariety. If $\bm{B}\in\mathsf{K}$ and $\bm{A}\leqslant\bm{B}$ is full and not epic in $\mathsf{K}$, then $\bm{A}\in\mathbb{S}(\mathsf{K}_{\textup{{rsi}}})$.
@@
$\textup{id}_{A}=(\textup{id}_{B}){\upharpoonright}_{A}=(\mathsf{Ker}(g)\cap\mathsf{Ker}(h)){\upharpoonright}_{A}=\mathsf{Ker}(g){\upharpoonright}_{A}\cap\mathsf{Ker}(h){\upharpoonright}_{A}=\mathsf{Ker}(g){\upharpoonright}_{A}.$
Therefore, $g{\upharpoonright}_{A}$ is injective as desired. As a consequence, $g{\upharpoonright}_{A}\colon\bm{A}\to\bm{C}$ is an embedding. Since $\bm{C}\in\mathsf{K}_{\textsc{rsi}}$, we conclude that $\bm{A}\in\mathbb{I}\mathbb{S}(\mathsf{K}_{\textup{{rsi}}})=\mathbb{S}\mathbb{I}(\mathsf{K}_{\textup{{rsi}}})=\mathbb{S}(\mathsf{K}_{\textup{{rsi}}})$ because $\mathsf{K}_{\textup{{rsi}}}$ is closed under isomorphisms. ∎
@@
Corollary 3.11 cannot be strengthened by concluding that also $\bm{B}\in\mathbb{S}(\mathsf{K}_{\textup{{rsi}}})$, nor even that $\bm{B}\in\mathbb{S}(\mathsf{K}_{\textup{{rfsi}}})$. To prove this, let $\bm{B}$ be the semilattice with meet-order $10<a<$ 1 and $\bm{A}\leqslant\bm{B}$ its subalgebra with universe $\{0,1\}$. While $\bm{A}\leqslant\bm{B}$ is easily seen to be full in the variety $\mathsf{SL}$ of semilattices, it is not epic because $\mathsf{SL}$ has the weak ES property (see, e.g., [21, p. 99]). Nonetheless, $\bm{B}$ is not a subalgebra of an FSI semilattice because, up to isomorphism, the only FSI semilattice is $\bm{A}$. ∎
@@ 4. Quasivarieties with a near unanimity term
In the presence of a near unanimity term, the task of determining whether a quasivariety has the ES property is simplified by the following result [8, Thm. 6.4]: [^5] In the statement of [8, Thm. 6.4], the occurrences of $n-1$ in Theorem 4.2 are replaced by $n$. However, its proof yields the stronger version reported here (see [25]).
@@
ordered under the relation given by $\langle\phi_{1},\dots,\phi_{m-1}\rangle\leqslant\langle\eta_{1},\dots,\eta_{m-1}\rangle$ if and only if $\phi_{i}\subseteq\eta_{i}$ for every $i\leqslant m-1$. We will apply Zorn’s Lemma to obtain a maximal element of $X$. Clearly, $X$ contains $\langle\theta_{1},\dots,\theta_{m-1}\rangle$. Consider a nonempty chain $C$ in $X$. For each $i\leqslant m-1$ let $C_{i}$ be the projection of $C$ on the $i$-th coordinate. Observe that $C_{i}$ is a nonempty chain in $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ because $C$ is a nonempty chain in $X$. We will prove that $\langle\bigcup C_{1},\dots,\bigcup C_{m-1}\rangle$ is an upper bound of $C$ in $X$. Proposition 3.6 implies that each $\bigcup C_{i}$ is a $\mathsf{K}$-congruence of $\bm{A}$. Furthermore, as $\theta_{i}$ is contained in every member of $C_{i}$ and $C_{i}$ is nonempty, we have $\theta_{i}\subseteq\bigcup C_{i}$ for every $i\leqslant m-1$. Lastly, observe that
$\left(\bigcup C_{1}\right)\cap\dots\cap\left(\bigcup C_{m-1}\right)=\bigcup\{\phi_{1}\cap\dots\cap\phi_{m-1}:\phi_{i}\in C_{i}\text{ for }i\leqslant m-1\}=\theta_{1}\cap\dots\cap\theta_{m-1},$ (1)
@@
where the the first inclusion holds because $\phi_{i}\in C_{i}$ for every $i\leqslant m-1$, and the last equality holds because $\langle\phi_{1}^{\prime},\dots,\phi_{m-1}^{\prime}\rangle\in C$. Hence, $\phi_{1}\cap\dots\cap\phi_{m-1}=\theta$ because $\theta=\theta_{1}\cap\dots\cap\theta_{m-1}$. Therefore, (1) holds, and so $\langle\bigcup C_{1},\dots,\bigcup C_{m-1}\rangle$ is an upper bound of $C$ in $X$. By Zorn’s Lemma the poset $X$ has a maximal element $\langle\phi_{1},\dots,\phi_{m-1}\rangle$. In particular, $\theta=\phi_{1}\cap\dots\cap\phi_{m-1}$.
@@
As $\theta$ is $m$-irreducible and $\theta\neq\phi_{1}\cap\dots\cap\phi_{k-1}\cap\phi_{k+1}\cap\dots\cap\phi_{m-1}$ for every $k\leqslant m-1$, we obtain
@@
$\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),b\rangle\in\theta_{j}\quad\text{forevery }j\leqslant n.$
To this end, consider $j\leqslant n$. As $\langle a_{i},b\rangle\in\phi_{i}\subseteq\theta_{j}$ for every $i\leqslant n$ such that $i\neq j$, we obtain $\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),\varphi^{\bm{B}}(b,\dots,b,a_{j},b,\dots,b)\rangle\in\theta_{j}$. Furthermore, since $\varphi$ is a near unanimity term, we have $\varphi^{\bm{B}}(b,\dots,b,a_{j},b,\dots,b)=b$. Hence, $\langle\varphi^{\bm{B}}(a_{1},\dots,a_{n}),b\rangle\in\theta_{j}$. This establishes the above display. Together with the assumption that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n}$, this implies $b=\varphi^{\bm{B}}(a_{1},\dots,a_{n})$. As $a_{1},\dots,a_{n}\in A$ and $\bm{A}\leqslant\bm{B}$, we conclude that $b\in A$, which is false. Hence, $\text{id}_{B}$ is $n$-irreducible as desired.
@@
The implication from left to right holds by Proposition 3.2. We will prove the other implication by contraposition. Suppose that $\mathsf{K}$ lacks the weak ES property. By Corollary 3.8 there exist $\bm{B}\in\mathsf{K}$ finitely generated and $\bm{A}\leqslant\bm{B}$ fully epic in $\mathsf{K}$. In view of Proposition 4.6, we may assume that $\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n-1}$ is a subdirect product for some $\bm{B}_{1},\dots,\bm{B}_{n-1}\in\mathsf{K}_{\textsc{rfsi}}$. ∎
@@
Let $\mathsf{K}$ be a class of lattices of bounded length such that $\mathbb{V}(\mathsf{K})$ is nontrivial. By Theorem 2.6, the FSI members of $\mathbb{V}(\mathsf{K})$ belong to $\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})$. Since the members of $\mathsf{K}$ have length $\leqslant n$ for some $n\in\mathbb{Z}^{+}$, so do those of $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K})$ by Łoś’ Theorem [7, Thm. V.2.9]. As the class operators $\mathbb{H}$ and $\mathbb{S}$ do not increase the cardinality of chains, we conclude that the FSI members of $\mathbb{V}(\mathsf{K})$ have length $\leqslant n$. Now, as $\mathbb{V}(\mathsf{K})$ is nontrivial, it has at least one SI member. Therefore, there exists $\bm{A}_{1}\in\mathbb{V}(\mathsf{K})_{\textsc{si}}$ of maximal length. Let $\textup{Cg}^{\bm{A}_{1}}(a,b)$ be the monolith of $\bm{A}_{1}$. As $a\neq b$, we may assume that $a<b$. Moreover, let $\bm{A}_{2}$ be the two-element chain $0<1$ viewed as a lattice, which is also SI. Consider the subalgebra $\bm{A}$ of $\bm{A}_{1}\times\bm{A}_{2}$ with universe $(A_{1}\times\{0\})\cup({\uparrow}b\times\{1\})$, where ${\uparrow}b\coloneqq\{c\in A_{1}:b\leqslant c\}$, and let $\bm{B}\coloneqq\textup{Sg}^{\bm{A}_{1}\times\bm{A}_{2}}(A\cup\{\langle a,1\rangle\})$. Clearly, $\bm{B}\leqslant\bm{A}_{1}\times\bm{A}_{2}$ is a subdirect product. We will prove that $\bm{A}\leqslant\bm{B}$ is fully epic in $\mathbb{V}(\mathsf{K})$, thus showing that $\mathbb{V}(\mathsf{K})$ lacks the weak ES property.